OpenStudy (anonymous):

How do I find the exact value of this trigonometric function? > cos 23π/6

OpenStudy (anonymous):

\[\cos(\frac{ 23\pi }{ 6 })\] \[=\cos(\frac{ 23(180) }{ 6 })\] \[=\cos(23\times30)\] \[=\cos(690-360)\] \[=\cos(330)\]

OpenStudy (campbell_st):

lots of ways of doing this 23/6 = 3 5/6 so its a 4th quadrant angle which has the same value as \[\cos(\pi/6)\] 23/6 = 24/6 - 1/6 you can also use the difference of 2 angles then expand \[\cos(\frac{24\pi}{6} - \frac{\pi}{6})\]

OpenStudy (anonymous):

If you plug in \[cos^{2}(330)\] into your calculator you will get \[\frac{ 3 }{ 4 }\] Then you can see that the exact value would be the square root of \[\frac{ 3 }{ 4 }\]